# Directly Constructing Low-Dimensional Solution Subspaces in Deep Neural Networks

### v1

@@ 3.5 Experimental Implementation Details
Table 1: Hyperparameter Configuration. Specific settings for the Subspace Construction phase (Frozen Backbone). Note the use of higher learning rates for Transformers to adapt the linear head to the fixed manifold.
Model Optimizer LR (Head) Weight Decay Epochs Batch Size
@@ 4.1 Main Results
Table 2 summarizes the classification accuracy across all three tasks. We report the raw accuracy and the relative difference ($\Delta$) compared to the full-dimensional Frozen Baseline.
@@
Table 2: Subspace Classification Results. Comparison of Full Fine-Tuning, Frozen Baseline (Full Dim), and JLL Subspace Projections. $\Delta$ denotes the performance difference relative to the Frozen Baseline. Note the stability of accuracy even at high compression rates (e.g., 12x).
Model / Task Method Dim ($k$) Ratio Accuracy $\Delta$ Baseline
ResNet-50 Full Fine-Tuning 2048 1x 79.95% -2.45%
(CIFAR-100) Frozen Baseline 2048 1x 82.40%
JLL Projection 1024 2x 82.29% -0.11%
JLL Projection 512 4x 81.96% -0.44%
JLL Projection 256 8x 81.69% -0.71%
JLL Projection 128 16x 81.19% -1.21%
ViT-B/16 Full Fine-Tuning 768 1x 93.84% -0.18%
(ImageNet-100) Frozen Baseline 768 1x 94.02%
JLL Projection 512 1.5x 93.88% -0.14%
JLL Projection 256 3x 94.32% +0.30%
JLL Projection 128 6x 93.90% -0.12%
JLL Projection 64 12x 93.86% -0.16%
BERT-base Full Fine-Tuning 768 1x 83.69% -0.05%
(MNLI) Frozen Baseline 768 1x 83.74%
JLL Projection 512 1.5x 83.60% -0.14%
JLL Projection 256 3x 83.58% -0.16%
JLL Projection 128 6x 83.36% -0.38%
JLL Projection 64 12x 83.70% -0.04%
4.2 Analysis of Results
@@
Robustness of the Solution Subspace: Across all architectures, the performance drop is minimal even under aggressive compression. For BERT (and ViT), reducing the dimension from 768 to 64 (a 12x reduction) results in a negligible accuracy loss. Similarly, ResNet-50 retains over 98.5% of its baseline performance when projected down to 128 dimensions. This confirms our hypothesis that the intrinsic solution manifold is significantly smaller than the ambient space allocated by the architecture.
5 Discussion and Conclusion
5.1 Theoretical Implications: The Flat Solution Manifold
@@
Our findings provide constructive evidence for the "Neural Flattening" hypothesis. While prior work measured the intrinsic dimension via local gradients or manifold estimation [24], our use of a global, data-independent linear projection offers a stronger guarantee: the solution manifold is not only low-dimensional locally but is also globally "well-behaved" (flat and linearly separable).
@@
The logic is as follows: If the class manifolds were highly curved or entangled in the high-dimensional space, a random linear projection would inevitably lead to significant class overlap and performance collapse (violating the margin requirements for JLL). The fact that this collapse does not occur implies that the backbone effectively linearizes the problem into a subspace $k\ll d$. The feature extractor acts as a "flattening engine," rendering the complex input data into a geometry so simple that even a random basis is sufficient to distinguish the classes.
@@ 5.2 Future Direction: Subspace-Native Distillation
Our experiments have empirically confirmed the existence of a robust solution subspace that captures nearly the entire performance of the teacher despite being a fraction of the size. This finding fundamentally challenges current Knowledge Distillation paradigms. Since we have proven that the relevant signal exists in the projected subspace $\tilde{h}=Rh$, we propose Subspace-Native Distillation. In this framework, the student is trained to directly construct the solution in the low-dimensional manifold defined by the fixed matrix $R$. The loss function shifts to:
@@
By targeting the intrinsic dimension directly, we hypothesize that the student can bypass the high-dimensional search problem entirely. The backbone of the student only needs enough capacity to map inputs to the $k$-dimensional target, potentially allowing for extreme parameter reduction while maintaining the teacher’s accuracy. This paves the way for a generation of "subspace-native" models that are optimized for the solution, not the search.
@@ 5.3 Conclusion
While high-dimensional width is crucial for solving the non-convex "search problem" during training, the final solution resides in a much simpler, low-rank subspace. This distinction provides the theoretical and empirical foundation for Subspace-Native Distillation. By shifting the distillation target from the noisy ambient space to this verified solution manifold, we can potentially train student models that are efficient by design rather than by approximation.

### v2

@@ 4.1 Implementation Details
Table 1: Hyperparameter Configuration. Settings for the linear probe (frozen backbone). These hyperparameters apply to both the full-dimensional baseline ($d$) and the low-dimensional subspace projections ($k$), ensuring a fair comparison. Note the use of higher learning rates (LR) for Transformers to adapt the linear head to the fixed manifold.
Model Optim. LR Wt. Decay Epochs Batch
@@
1. Vision (CNN): ResNet-50 on CIFAR-100 We utilize a ResNet-50 backbone and fine-tune it on CIFAR-100. The feature vector is extracted from the final pooling layer ($d=2048$) before the fully connected layer.
  • Preprocessing: We employ data augmentation to prevent overfitting on the small $32\times 32$ CIFAR images, including random crop, horizontal flip, color jitter, and cutout regularization (DeVries and Taylor, 2017) (1 hole, max size 8).
  • Optimization: The model is optimized using stochastic gradient descent (SGD) with momentum (0.9). For the Subspace phase, we train for 5 epochs with a learning rate of $1e^{-2}$ and weight decay of $5e^{-4}$.
2. NLP (Transformer): BERT-base on MNLI We employ the bert-base-uncased model ($d=768$) and fine-tune it on MNLI. The input consists of premise-hypothesis pairs with a maximum sequence length of 128 tokens.
  • Feature Extraction: We utilize the pooler_output (the embedding of the [CLS] token processed by a dense layer and Tanh activation) as the input $h$ for our projections.
  • Optimization: We use the AdamW optimizer (Loshchilov and Hutter, 2019). While full fine-tuning requires a conservative learning rate ($2e^{-5}$), the subspace classification phase allows for a higher learning rate ($1e^{-3}$) to rapidly converge the linear head on the frozen features.
3. Vision (Transformer): ViT-B/16 on ImageNet-100 We utilize a Vision Transformer (ViT-B/16) and fine-tune it on ImageNet-100.
  • Preprocessing: Images are resized to $224\times 224$. Similar to ResNet, we apply cutout (max. size 32) during training to encourage robust feature learning. The [CLS] token of the last hidden state serves as the feature vector.
  • Optimization: Training is performed with AdamW. As with BERT, we utilize a learning rate of $1e^{-3}$ for the subspace classifiers to effectively regress the target decision boundaries within 3 epochs.
Dimensionality Reduction Factors. We vary the target dimension $k$ across a wide spectrum. For ResNet-50 ($d=2048$), we evaluate subspaces of width $k\in\{1024,512,256,128\}$, corresponding to compression factors up to $16\times$. For BERT and ViT ($d=768$), we evaluate $k\in\{512,256,128,64\}$, reaching compression factors of $12\times$.
@@ 5.1 Main Results
Table 2 summarizes the classification accuracy across all three tasks. We report the raw accuracy and the relative difference compared to the full-dimensional fine-tuned baseline.
@@
Table 2: Subspace Classification Results. Comparison of Fine-tuned Baseline (full dimension), and JL subspace projections. $\Delta$ denotes the performance difference relative to the Baseline. Note the stable accuracy even at high compression rates (e.g., $12\times$).
Method Dim ($k$) Ratio Acc. $\Delta$ Base
ResNet-50 (CIFAR-100)
Fine-tuned Baseline 2048 $1\times$ 82.40%
JL Projection 1024 $2\times$ 81.32% -1.08%
JL Projection 512 $4\times$ 81.20% -1.20%
JL Projection 256 $8\times$ 80.89% -1.51%
JL Projection 128 $16\times$ 80.19% -2.21%
ViT-B/16 (ImageNet-100)
Fine-tuned Baseline 768 $1\times$ 94.02%
JL Projection 512 $1.5\times$ 93.52% -0.50%
JL Projection 256 $3\times$ 93.98% -0.04%
JL Projection 128 $6\times$ 93.90% -0.12%
JL Projection 64 $12\times$ 93.28% -0.74%
BERT-base (MNLI)
Fine-tuned Baseline 768 $1\times$ 83.74%
JL Projection 512 $1.5\times$ 83.93% +0.19%
JL Projection 256 $3\times$ 83.79% +0.05%
JL Projection 128 $6\times$ 83.79% +0.05%
JL Projection 64 $12\times$ 83.64% -0.10%
@@
Robustness of the Solution Subspace: Across all architectures, the performance drop is minimal even under aggressive compression. For BERT (and ViT), reducing the dimension from 768 to 64 (a $12\times$ reduction) results in less than $1\%$ accuracy drop. Similarly, ResNet-50 retains over $97\%$ of its baseline performance when projected down to 128 dimensions. This confirms our hypothesis that the intrinsic solution manifold is significantly smaller than the ambient space allocated by the architecture. In fact, for BERT, mild compression ($k=512$) yields a slight accuracy gain ($0.19\%$), implying that the ambient representation contains noise components that the projection filters out.
To provide a qualitative intuition, we visualize the feature space geometry using $t$-SNE in Figure 1. Figure 1(a) displays the raw, high-dimensional representations ($d=768$) of the fully fine-tuned model, exhibiting clear class separation. Figure 1(b) shows the same samples projected into the random low-dimensional subspace ($k=128$) via the fixed JL matrix.
Crucially, the cluster structure remains intact for the JL projection. The distinct separation between classes is preserved, and the local neighborhood relationships appear undisturbed. This supports our hypothesis that the essential geometry required for classification resides in a subspace that survives data-independent random contraction.
(a) Raw Features (768D)
(b) JL Projection (128D)
Figure 1: $t$-SNE visualization of the same examples before and after a frozen JL projection. The same seed and the same $t$-SNE hyperparameters have been used across all visualizations.
6 Ablation
To validate that the solution geometry is intrinsic to the data manifold rather than an artifact of specific projection techniques, we compare our oblivious JL approach against two distinct and data-dependent baselines:
PCA projection. We compute the principal components of the frozen training features and project the data onto the top-$k$ components. This represents the theoretically optimal linear projection for maximizing variance preservation.
Learned projection. Instead of a fixed random matrix $R$, we treat the projection matrix as a trainable parameter, which is initialized randomly to the same values as in the JL projection. This is updated end-to-end alongside the classifier head via backpropagation.
Results. Table 3 presents the comparative analysis across the architectures. We reproduce the JL results from Table 2 for direct comparison. As expected, PCA generally yields the highest accuracy because it is data-dependent, such that it aligns the projection basis with the axes of maximum information. However, the oblivious JL projection remains competitive to both ablations, the PCA projection, as well as the learned projection. This parity validates the ’Heavy-Tailed’ spectral hypothesis Martin and Mahoney (2021): it implies that the discriminative signal is so structurally dominant over the noise bulk that precise eigen-alignment (via PCA) is redundant. The solution geometry is sufficiently robust to survive blind, random contraction.
Vision (ResNet-50). While PCA is the upper bound, JL projections remain on par. Notably, at $k=512$, JL achieves 81.20% compared to 80.56% for the learned matrix. Since the learned matrix was initialized from the exact same JL values, this performance drop suggests that optimizing the basis on the training set led to overfitting, whereas the frozen random features preserved the global geometry robustly. Transformers (ViT & BERT). In the Transformer domain, we observe a similar pattern. Notably, for ViT at $k=256$, JL outperforms both PCA and learned projections.
We further compare the manifold structure of the subspaces in Figure 2. While $t$-SNE introduces inherent non-linear distortions to the cluster shapes, the topological structure produced by the PCA projection (Figure 2(a)) and the learned projection (Figure 2(b)) is qualitatively comparable to our oblivious JL projection (Figure 1(b)).
Table 3: Ablation Study. Comparison of JL projections (ours) against data-dependent (PCA) and optimization-dependent (Learned) projections. JL projections achieve parity with data-dependent methods, confirming the robustness of the solution geometry.
Dim ($k$) JL (Ours) PCA Learned
ResNet-50 (CIFAR-100)
1024 ($2\times$) 81.32% 81.42% 81.31%
512 ($4\times$) 81.20% 81.30% 80.56%
256 ($8\times$) 80.89% 81.35% 80.96%
128 ($16\times$) 80.19% 81.21% 80.51%
ViT-B/16 (ImageNet-100)
512 ($1.5\times$) 93.52% 93.92% 93.74%
256 ($3\times$) 93.98% 93.90% 93.78%
128 ($6\times$) 93.90% 94.06% 93.74%
64 ($12\times$) 93.28% 93.98% 93.42%
BERT-base (MNLI)
512 ($1.5\times$) 83.93% 83.96% 83.83%
256 ($3\times$) 83.79% 83.83% 83.88%
128 ($6\times$) 83.79% 83.70% 83.72%
64 ($12\times$) 83.64% 83.70% 83.98%
(a) PCA Projection (128D)
(b) Learned Projection (128D)
Figure 2: $t$-SNE visualization of the same examples for PCA projection and learned projection. The same seed and the same $t$-SNE hyperparameters have been used across all visualizations.
7 Discussion
7.1 Implications: The Flat Solution Manifold
@@
Our findings provide constructive evidence for the Neural Flattening hypothesis. While prior work measured the intrinsic dimension via local gradients or manifold estimation (Ansuini et al., 2019), our use of a global, data-independent linear projection offers a stronger guarantee: the solution manifold is not only low-dimensional locally but is also globally well-behaved (flat and linearly separable).
@@
The logic is as follows: If the class manifolds were highly curved or entangled in the high-dimensional space, a random linear projection would inevitably lead to significant class overlap and performance collapse (violating the margin requirements for JLL). The fact that this collapse does not occur implies that the backbone effectively linearizes the problem into a subspace $k\ll d$. The feature extractor acts as a flattening mechanism, rendering the complex input data into a geometry so simple that even a random basis is sufficient to distinguish the classes.
@@ 7.2 Future Direction: Subspace-Native Distillation
Our experiments have empirically confirmed the existence of a robust solution subspace that captures nearly the entire performance of the teacher despite being a fraction of the size. This finding challenges current knowledge distillation paradigms. Since we have shown that the relevant signal exists in the projected subspace $\tilde{h}\propto Rh$, the practically relevant next step—which we leave for future work—is to exploit this for training and inference. For that goal, we propose the idea of Subspace-Native Distillation. In this framework, the student is trained to directly construct the solution in the low-dimensional manifold defined by the fixed matrix $R$. The loss function shifts to:
@@
By targeting the intrinsic dimension directly, we hypothesize that the student can bypass the high-dimensional search problem entirely. The backbone of the student only needs enough capacity to map inputs to the $k$-dimensional target, potentially allowing for extreme parameter reduction while maintaining the teacher’s accuracy. This paves the way for a generation of subspace-native models that are optimized for the solution, not the search.
7.3 Limitations
Layer-wise Analysis. While our findings provide evidence for the disconnect between optimization width and solution geometry, our analysis focuses exclusively on the final latent representation. As established by Ansuini et al. (2019), the intrinsic dimension of deep networks is not uniform across layers; it typically peaks in intermediate layers before collapsing towards the output. It remains an open question whether our “optimization scaffolding” hypothesis implies that intermediate layers must remain high-dimensional to facilitate gradient flow and feature disentanglement, or if the solution manifold is already constructible at earlier stages.
Task Generalization.
Furthermore, our evaluation is centered on discriminative classification tasks where the number of classes is small relative to the model width ($C<d$). In this regime, theoretical frameworks like Neural Collapse predict low-rank separability. However, tasks with continuous output spaces (e.g., Variational Autoencoders, diffusion models) or massive output vocabularies (where $C\gg d$, such as LLMs) may inherently require higher-dimensional manifolds to preserve information that oblivious random projections might obliterate.
@@ 7.4 Conclusion
While high-dimensional width is crucial for solving the non-convex optimization problem during training, the final solution resides in a much simpler, low-rank subspace. This distinction provides the theoretical and empirical foundation for Subspace-Native Distillation. By shifting the distillation target from the noisy ambient space to this valid solution manifold, we can potentially train student models that are efficient by design rather than by approximation.